Use the quadratic formula to determine the x-intercepts (if any) of the following function. Then evaluate the function for several values of x. I need all possible answers and the work this was so confusing.
f(x) = 7x^2 - 2

Answer by dabanfield(803)

f(x) = 7x^2 - 2
X-intercepts are values of x where:
0 = 7x^2 - 2
Rewrite this as:
7x^2 + 0*x - 2 = 0
In this standard quadratic form a=7, b=0 and c=-2.
Substitute these values for a, b and c in the quadratic formula.
To get values for f(x) for different values of x just substitute for x in the expression 7x^2 - 2. For example:
f(1) = 7*1^2 - 2
f(1) = 7*1 - 2
f(10 = 7 - 2 = 5
f(-1) = 7*(-1)^2 - 2
f(-1) = 7*1 - 2
f(-1) = 7 - 2 = 5
f(2) = 7*2^2 - 2
f(2) = 7*4 - 2
f(2) = 28 - 2 = 26